Write an equation in point-slope form for the line through the point (-2,1) parallel to the line with equation 3x - y = -9
Answer by mananth(16946)

Write an equation in point-slope form for the line through the point (-2,1) parallel to the line with equation 3x - y = -9
3x-y = -9
y=3x+9
compare with equation y= mx+b
m=3 = slope
a parallel line will have same slope of m=3
point = (-2,1)
general equation is
y-y1=m(x-x1)
where x,1y1 are the co ordinates of the point
y-1=3(x-(-2))
y-2=3x+6
y=3x+6+2
y=3x+8